# Win Ratio with Multiple Thresholds for Composite Endpoints

### v4

@@ 2.2 Win ratio with multiple thresholds
Figure 1: A schematic procedure of pairwise comparisons between two participants, $i$ and $j$, in the WR-MT with thresholds $(d,0)$ for the survival time endpoint and $(t_{1},0)$ for the time-to-hospitalization endpoint.
To formally introduce our approach, we first define the comparison functions for survival time and time-to-hospitalization endpoints, under thresholds $d\geq 0$ and $t_{1}\geq 0$, as $\mathscr{D}_{ij}(d)$ and $\mathscr{T}_{ij1}(t_{1})$, as follows:
@@
Once the win-loss score is computed for each pair, the estimation of the win ratio and relevant hypothesis testing can proceed following the general procedure outlined in Section 2.1. Hence, the corresponding 4-stage WR-MT can be denoted as
$\text{WR-MT}(d,t_{1},0,0)=\frac{\sum_{i=1}^{n}Z_{i}\sum_{j=1}^{n}(1-Z_{j})I\left\{U_{ij}(d,t_{1},0,0)=1\right\}}{\sum_{i=1}^{n}Z_{i}\sum_{j=1}^{n}(1-Z_{j})I\left\{U_{ij}(d,t_{1},0,0)=-1\right\}}.$ (1)
In general, the number of stages of WR-MT can be increased but should be pre-specified in the study design. For example, a 6-stage WR-MT can be formed by specifying a sequence $\{d_{1},t_{11},d_{2},t_{12},0,0\}$ and let the comparison go along the thresholds sequentially. To this end, the thresholds for the same information in later stages should always be smaller than the ones in higher stages (e.g., $d_{2}<d_{1},t_{12}<t_{11}$) to make the added stage non-trivial. In practice with two layers, we recommend a 4-stage WR-MT, which usually provides sufficient control of the priorities between survival time and time-to-hospitalization endpoints. In addition, we recommend using zero as the threshold for the last two stages, i.e., the steps considered in standard WR. This is aligned with the recommendation by Redfors et al. (2020) for the continuous endpoints to reduce ties to the maximum extent given that the increased number of ties can reduce the power. In Section 3.3, we will show in simulation that by employing zero in the last two stages, the proposed method will leave the same number of ties as in WR. We also note that there are situations where a minimum non-zero threshold may be employed for the comparison to exclude the clinical indifference area. For example, sometimes a difference in the Kansas City Cardiomyopathy Questionnaire Overall Score of 5 points is used as clinically significant (Voors et al., 2022). In that case, the final stages may keep non-zero thresholds.
In the 4-stage WR-MT illustrated above, there are two opportunities for the time-to-hospitalization endpoint to contribute to the result of a pairwise comparison, as opposed to the single opportunity provided by WR. Such additional opportunity releases the strict hierarchy. As two extreme examples, when $d=+\infty,t_{1}=0$, WR-MT prioritizes the time-to-hospitalization endpoint strictly like WR prioritizes the survival time endpoint. In contrast, when $d=0,t_{1}=0$, WR-MT reduces to standard WR. Therefore, by varying the thresholds, one may control the priority between two endpoints from one end to the other. Such flexibility allows the convenient inclusion of clinical priors.
@@ 2.3 Adaptive thresholds
In this section, we introduce a pre-specified data-driven adaptive approach to choose the thresholds in the absence of sufficient clinical information. This adaptive approach is based on the quantile of survival time and time-to-hospitalization differences among all pairwise comparisons, and can help select the thresholds in an intuitive way. To begin with, we define two sequences of all non-zero differences in pairwise comparisons for survival time and time to hospitalization accordingly:
@@
These two sequences take comparisons between two participants from the same (i.e., both receive treatment or control) or contrasting groups (i.e., one receives treatment and the other receives control) into account, which is essential for gaining the adaptive property (will be discussed shortly). Then, with the calipers $c_{1},c_{2}\in[0,1]$ specified, we calculate the empirical $c_{1},c_{2}$ quantile values of these two sequences correspondingly, denoted as $q_{D}^{c_{1}}=\hat{q}_{c_{1}}(\mathbb{D})$ and $q_{1}^{c_{2}}=\hat{q}_{c_{2}}(\mathbb{T}_{1})$. Next, we set $(d,t_{1})=(q_{D}^{c_{1}},q_{1}^{c_{2}})$, where, loosely speaking, the relationship between $c_{1}$ and $c_{2}$ controls the weights of survival time and time-to-hospitalization endpoints. Intuitively, a larger caliper $c$ corresponds to a larger threshold, indicating that less attention is paid to an endpoint. Therefore, $c_{1}\leq c_{2}$ is usually adopted as we shall still prioritize the higher layer even though the strict hierarchy is intended to be released. The influence of calipers on the weight between endpoints is more predictable and intuitive than that of thresholds themselves since the effect of thresholds relies on their relative scale with the observed differences. This helps the interpretation of tuning parameters.
@@
To control the weight more straightforwardly and equivalently, we consider employing $c_{1}=c_{2}=c$ and a weight parameter $w>0$: after obtaining $q_{D}^{c}=\hat{q}_{c}(\mathbb{D})$ and $q_{1}^{c}=\hat{q}_{c}(\mathbb{T}_{1})$, we set $(d,t_{1})=(q_{D}^{c},q_{1}^{c}/w)$, where the value of $w$, instead of the relationship between $c_{1}$ and $c_{2}$ controls the weight between two endpoints. A smaller $w$ indicates that less attention is paid to the time-to-hospitalization endpoint, as it increases the threshold for this endpoint. Since the remaining tuning parameters are $w,c$, we denote the resulted $\text{WR-MT}(q_{D}^{c},q_{1}^{c}/w,0,0)$ as $\text{WR-AT}(w)^{c}$, which stands for the WR with adaptive thresholds by weight $w$ and shared caliper $c$. In Appendix A, we present the generalized formulas for adaptive thresholds under multiple non-fatal events for enhanced applicability.
@@
In the process of constructing adaptive thresholds, we note that only non-zero differences are included in the sequences to exclude the results of self-comparison and ties. This ensures the subsequent quantiles are non-trivial (i.e., $d\neq 0,t_{1}\neq 0$). The caliper $c$ then indicates $100c\%$ smallest non-zero absolute differences shall be considered as ties in Stages 1 and 2. As an empirical rule of thumb, we recommend using $c=20\%$ as the default value when clinical priors are lacking, and its impact on WR-AT will be demonstrated in the simulation section. When the default caliper value is employed, for simplicity, we omit the superscript $c=20\%$ in the above notations and denote $q_{D}=q_{D}^{20\%}=\hat{q}_{20\%}(\mathbb{D}),q_{1}=q_{1}^{20\%}=\hat{q}_{20\%}(\mathbb{T}_{1}),\text{WR-AT}(w)=\text{WR-AT}(w)^{20\%}$. As for the weight parameter $w$, a smaller $w$ assigns a smaller weight to the time-to-hospitalization endpoint by elevating the thresholds for differences in the time to hospitalization in Stage 2. Since the differential clinical importance between two endpoints is already addressed by the hierarchy and the caliper, the default value $w=1$ is often an adequate choice without external clinical knowledge. In addition, we will show that WR-AT is generally robust to the choice of several common values of $w$ through simulations. Operationally, the adaptive thresholds in $\text{WR-AT}(w)^{20\%}$ can be obtained as follows:
@@
We offer additional remarks before presenting the simulation study. First, the validity of the hypothesis test based on WR-MT remains unaffected by the specification of thresholds. Under the null hypothesis of no treatment effect across endpoints, the choice of thresholds is not expected to introduce spurious signals or differential outcomes, thereby preserving the nominal type I error rate asymptotically. This property provides some justifications for using adaptive thresholds derived empirically from the observed data (see Table 2 for empirical type I error rates for WR-AT). However, as elaborated below, the choice of thresholds can influence power. In this regard, threshold specification under WR-MT is reminiscent of test statistic specification in permutation tests, which has an impact on power without compromising validity. Second, the use of adaptive thresholds leads to more adaptive win ratio analysis. For example, when the treatment only has an effect on extending the time to hospitalization, WR-AT improves power compared to the standard WR. On the other hand, when the treatment only has an effect on survival, WR-AT maintains comparable power to standard WR. This adaptive property arises from estimating thresholds using pooled comparisons (i.e., pairwise comparisons among all participants, regardless of treatment group) and then applying these thresholds to comparisons between treatment arms. If no treatment effect exists on the survival time layer, the distribution of absolute survival differences between contrasting treatment groups resembles that of the pooled population. In this case, using the default $c=20\%$, approximately 20% of non-zero differences between treatment and control pairs will result in ties and move the comparison to the next endpoint (e.g., time to hospitalization). This reduces the influence of random variation on the survival layer. Conversely, when a treatment effect exists on survival time, the distribution of absolute survival differences between treatment groups will shift rightward relative to the pooled distribution. As a result, the same 20% quantile threshold now classifies fewer contrasting pairs as ties in Stage 1, thereby allowing more comparisons to be resolved at the survival layer. Taken together, these properties make WR potentially appealing in practice, especially when specifying fixed thresholds a priori is challenging.
@@ 3 Simulation Study
In this section, we examine the performance of WR-AT empirically via simulation. We first compare WR-AT with standard WR regarding empirical power (after validating their ability to control type I error), and then analyze the influence that caliper $c$ and weight $w$ have on the empirical power achieved by WR-AT. Although the proposed WR-MT method can accommodate multiple nonfatal events, without loss of generality, we focus on the simpler setting with two endpoints (survival time and time to hospitalization).
@@ 3.1 Simulation setup
We consider a two-arm clinical trial with a total sample size $n=2000$ and equal allocation, where half of them are assigned to the treatment group $Z=1$ and the other half to the control group $Z=0$. We simulate the two correlated times to represent potentially latent survival time in days and time to hospitalization in days using a copula approach. Specifically, following Luo et al. (2015), we adopt the Gumbel-Hougaard copula with a bivariate distribution that has exponential margins. We let $h_{D}(Z)=\lambda_{D}\exp(-\alpha_{D}Z)$ be the hazard rate for death event, and $h_{T_{1}}(Z)=\lambda_{T_{1}}\exp(-\alpha_{T_{1}}Z)$ be the hazard rate for hospitalization event. Then, the vector of survival time and time to hospitalization (in days) $(D^{*},T_{1}^{*})$ has the joint survival function:
@@
where $\beta\geq 1$ controls the correlation between the two endpoints (i.e., Kendall’s concordance: $1-1/\beta$). We fix parameters $\lambda_{D}=0.0008,\lambda_{T_{1}}=0.0022$ and let $\alpha_{D},\alpha_{T_{1}}\in\{0,0.1,0.2,0.3\}$, in which $\{0,0.1,0.2,0.3\}$ stand for no, very weak, weak, and modest treatment effects. We assume treatment is not worse than control (in terms of extending survival time or time to hospitalization). In our simulation, it is possible to have sample $i$ with $T_{1i}^{*}\geq D^{*}_{i}$, which indicates there is no hospitalization event for $i$. Observed survival time and time to hospitalization are then obtained by performing administrative censoring after the scheduled days of follow-up for all participants, with the scheduled follow-up length denoted as FU. A significance level of $\alpha=0.05$ for a two-sided test is employed throughout.
We compare WR-AT (with default weight and caliper choices, i.e., $\text{WR-AT}(w=1)^{c=20\%}$) with standard WR and examine the performance of WR-AT under different choices of weights and calipers. Three key factors, follow-up time, correlation between endpoints, and treatment effect size, are varied and detailed in Table 1 for simulation scenarios S1-S8. Simulation results under additional scenarios are presented in Appendix B. The empirical power is based on 2000 simulation replicates, and the empirical type I error rate calculation is based on 5000 replicates.
All computations are implemented in R, and include function calls from gumbel (Dutang, 2018), doParallel (Microsoft Corporation and Weston, 2022), dplyr (Wickham et al., 2021), ggplot2 (Wickham, 2016). An R package “WRMT” that implements the proposed method has been developed and is available at https://github.com/YH-Mou/WRMT.
@@ 3.2 Performance of WR-AT
The empirical power of WR and WR-AT under simulation scenarios S1-S8 (i.e., under the non-null scenarios) is summarized in Figure 2. To begin with, in simulation scenarios S1 and S2, WR-AT continuously increases the power from standard WR by allowing the information on the time-to-hospitalization layer to contribute to the inference about the treatment effect. The difference between WR-AT and WR is more apparent with a longer follow-up time. Since longer follow-up time indicates more observed death events on the survival time layer, which brings more random differences, WR-AT increases the power by more frequently leveraging the time-to-hospitalization layer and bypassing the random difference on the survival time layer. With a shorter follow-up time, WR maintains good performance as lower death rates allow a higher contribution from the time-to-hospitalization layer to pairwise comparison results. In simulation scenarios S3 and S4, the treatment effect is on the survival time layer with random differences on the time-to-hospitalization layer. Although more frequently exploiting the time-to-hospitalization layer is expected to reduce power, WR-AT presents comparable power to WR in simulation scenario S4, which is a result of employing the proposed adaptive thresholds. In simulation scenario S3, however, it is noticeable that the added correlation between endpoints lowers the power of both methods and WR-AT appears to be less powerful than WR. This phenomenon will be further investigated in Section 3.3.
In simulation scenarios S1-S4, we consider including only random differences on one of the endpoints to demonstrate the performance of WR-AT under extreme situations. In practice, however, if prior information indicates no treatment effect on an endpoint, including it in the hierarchical structure should be discouraged when using either WR-AT or standard WR. In simulation scenarios S5-S8, we consider more common settings where there are treatment effects on both endpoints but with different magnitudes. In simulation scenarios S5 and S6, where the treatment effect on the survival time endpoint is lower than that on the time-to-hospitalization endpoint, WR-AT becomes more powerful than WR. Similar to simulation scenarios S1 and S2, a longer follow-up time may reduce power as more pairwise comparison results are determined by the survival time endpoint, for which only the treatment effect is weak. WR-AT appears to be more robust in terms of maintaining power in such a scenario. Different from the trend observed in simulation scenarios S1 and S2, the power trend of both WR and WR-AT here first increases from 250-day follow-up to 500-day follow-up and then starts to decrease, which is likely a result of the difference between having very weak and completely no treatment effects on the survival time layer. Additionally, in simulation scenarios S7 and S8 where the treatment effect on the time-to-hospitalization endpoint is weak, WR-AT achieves comparable power with WR (similar to simulation scenario S4). Finally, the empirical type I error in simulation scenario S0 with Kendall’s concordance set to 0 and 0.5 is presented in Table 2. At the nominal level of $\alpha=0.05$, both WR and WR-AT adequately control empirical type I error rates, with observed values falling within the expected range of Monte Carlo variation.
@@
Figure 2: Empirical power of WR-AT and WR tests in simulation scenario: (S1) modest treatment effect on the time-to-hospitalization endpoint only, assuming a correlation between survival time and time-to-hospitalization endpoints, (S2) modest treatment effect on the time-to-hospitalization endpoint only, assuming no correlation, (S3) modest treatment effect on the survival time endpoint only, assuming a correlation, (S4) modest treatment effect on the survival time endpoint only, assuming no correlation, (S5) very weak treatment effect on the survival time endpoint and weak treatment effect on the time-to-hospitalization endpoint, assuming a correlation, (S6) very weak treatment effect on the survival time endpoint and weak treatment effect on the time-to-hospitalization endpoint, assuming no correlation, (S7) weak treatment effect on the survival time endpoint and very weak treatment effect on the time-to-hospitalization endpoint, assuming correlation, (S8) weak treatment effect on the survival time endpoint and very weak treatment effect on the time-to-hospitalization endpoint, assuming no correlation.
@@
Table 2: Empirical type I error rates (%) obtained by WR-AT and WR under the null (simulation scenario S0) with Kendall’s concordance 0 and 0.5. The acceptable range of empirical type I error rate with Monte Carlo variation under 5000 replicates is 4.41% to 5.64% for the 5% nominal level.
Scenario Kendall’s Concordance FU WR WR-AT

### v5

@@ 2.2 Alternating multiple thresholds across endpoints
Figure 1: A schematic procedure of pairwise comparisons between two participants, $i$ and $j$, in the FS-MT test with thresholds $(d,0)$ for the survival time endpoint and $(t_{1},0)$ for the time-to-hospitalization endpoint.
To formally introduce our approach, we first define the comparison functions for survival time and time-to-hospitalization endpoints, under thresholds $d\geq 0$ and $t_{1}\geq 0$, denoted by $\mathscr{D}_{ij}(d)$ and $\mathscr{T}_{ij1}(t_{1})$, as follows:
@@
Once the win-loss score is computed for each pair, hypothesis testing with the FS test can proceed following the general procedure outlined in Section 2.1.
In general, the number of stages of a FS-MT test can be increased but should be pre-specified. For example, a 6-stage FS-MT test can be formed by specifying a sequence $\{d_{1},t_{11},d_{2},t_{12},0,0\}$ and by requiring the comparison to go along the thresholds sequentially. To this end, the thresholds for the same information in later stages should always be smaller than the ones in higher stages (e.g., $d_{2}<d_{1},t_{12}<t_{11}$) to make the added new stage non-trivial. In practice with two layers, we recommend a 4-stage FS-MT test, which usually provides sufficient control of the priorities between survival time and time-to-hospitalization endpoints. In Section 3.3, we will show in simulation that by employing the last two stages, which correspond exactly to those used in the standard FS test, the proposed method yields the same number of ties as the FS test.
In the 4-stage FS-MT test illustrated above, there are two opportunities for the time-to-hospitalization endpoint to contribute to the result of a pairwise comparison, as opposed to the single opportunity provided by the standard FS test. Such additional opportunity releases the strict hierarchy. As two extreme examples, when $d=+\infty,t_{1}=0$, the FS-MT test prioritizes the time-to-hospitalization endpoint strictly just like the standard FS test prioritizes the survival time endpoint. In contrast, when $d=0,t_{1}=0$, the FS-MT test reduces to standard FS test. Therefore, by varying the thresholds, one may control the priority between two endpoints from one end to the other. Such flexibility allows the convenient inclusion of clinical priors.
@@ 2.3 Adaptive thresholds
To improve the practical utility of the FS-MT testing procedure, in this section, we introduce a pre-specified data-driven adaptive approach to choose the thresholds. This can be useful in the absence of sufficient clinical information that determines the multiple thresholds. To begin with, we define two sequences of all non-zero differences in pairwise comparisons for survival time and time to hospitalization accordingly:
@@
These two sequences take comparisons between two participants from the same (i.e., both receive treatment or control) or contrasting groups (i.e., one receives treatment and the other receives control) into account, which is essential for gaining the adaptive property (will be discussed shortly). Then, with the calipers $c_{1},c_{2}\in[0,1]$ specified, we calculate the empirical $c_{1},c_{2}$ quantile values of these two sequences correspondingly, denoted as $q_{D}^{c_{1}}=\hat{q}_{c_{1}}(\mathbb{D})$ and $q_{1}^{c_{2}}=\hat{q}_{c_{2}}(\mathbb{T}_{1})$. Next, we set $(d,t_{1})=(q_{D}^{c_{1}},q_{1}^{c_{2}})$, where, loosely speaking, the relationship between $c_{1}$ and $c_{2}$ controls the weights of survival time and time-to-hospitalization endpoints. Intuitively, a larger caliper $c$ corresponds to a larger threshold, indicating that less emphasis is placed on an endpoint. Therefore, $c_{1}\leq c_{2}$ is usually adopted as we shall still prioritize the higher layer even though the strict hierarchy is intended to be released. In general, the influence of calipers on the weight between endpoints may be more intuitive than that of thresholds themselves since the thresholds rely on their relative scale with the observed differences.
@@
To control the weight in a practical fashion, we consider employing $c_{1}=c_{2}=c$ and a weight parameter $w>0$: after obtaining $q_{D}^{c}=\hat{q}_{c}(\mathbb{D})$ and $q_{1}^{c}=\hat{q}_{c}(\mathbb{T}_{1})$, we set $(d,t_{1})=(q_{D}^{c},q_{1}^{c}/w)$, where the value of $w$, instead of the relationship between $c_{1}$ and $c_{2}$ controls the weight between two endpoints. A smaller $w$ indicates that less emphasis is placed on the time-to-hospitalization endpoint, as it increases the threshold for this endpoint. Since the remaining tuning parameters are $w,c$, we denote the resulted $\text{FS-MT}(q_{D}^{c},q_{1}^{c}/w,0,0)$ test as $\text{FS-AT}(w)^{c}$ test, which stands for the FS test with adaptive thresholds by weight $w$ and shared caliper $c$. In Supporting Information S1, we present the generalized formulas for adaptive thresholds under multiple non-fatal events for completeness.
@@
In the process of constructing adaptive thresholds, we note that only non-zero differences are included in the sequences to exclude the results of self-comparison and ties. This ensures the subsequent quantiles are non-trivial (i.e., $d\neq 0,t_{1}\neq 0$). The caliper $c$ then indicates $100c\%$ smallest non-zero absolute differences shall be considered as ties in Stages 1 and 2. As an empirical rule of thumb, we recommend using $c=20\%$ as the default value when clinical priors are lacking, and its impact on the FS-AT test will be demonstrated in ensuing simulation study. When the default caliper value is employed, for simplicity, we omit the superscript $c=20\%$ in the above notations and denote $q_{D}=q_{D}^{20\%}=\hat{q}_{20\%}(\mathbb{D}),q_{1}=q_{1}^{20\%}=\hat{q}_{20\%}(\mathbb{T}_{1}),\text{FS-AT}(w)=\text{FS-AT}(w)^{20\%}$. As for the weight parameter $w$, a smaller $w$ assigns a smaller weight to the time-to-hospitalization endpoint by elevating the thresholds for differences in the time to hospitalization in Stage 2. Since the differential clinical importance between two endpoints is already addressed by the hierarchy and the caliper, the default value $w=1$ is often an adequate choice without external clinical knowledge. In addition, we will show that the FS-AT test is generally robust to the choice of several common values of $w$ through simulations.
Operationally, the adaptive thresholds in $\text{FS-AT}(w)^{20\%}$ test can be obtained as follows:
@@
We provide several additional remarks before presenting the simulation study. First, the validity of the hypothesis test with the FS-MT test remains unaffected by the specification of thresholds. Under the null hypothesis of no treatment effect across endpoints, the choice of thresholds is not expected to introduce spurious signals or differential outcomes, thereby preserving the nominal type I error rate asymptotically. This property provides some justifications for using adaptive thresholds derived empirically from the observed data (see Table 2 for empirical type I error rates for the FS-AT test). However, as elaborated below, the choice of thresholds can influence power. In this regard, threshold specification under the FS-MT test is reminiscent of test statistic specification in permutation tests, which has an impact on power without compromising the test validity. Second, the use of adaptive thresholds leads to more adaptive hypothesis testing procedure (but still based on a pre-specified rule). For example, when the treatment only has an effect on extending the time to hospitalization, the FS-AT test improves power compared to the standard FS test. On the other hand, when the treatment only has an effect on survival, the FS-AT test maintains comparable power to the standard FS test. This adaptive property arises from estimating thresholds using pooled comparisons (i.e., pairwise comparisons among all participants, regardless of treatment group) and then applying these thresholds to comparisons between treatment arms. If no treatment effect exists on the survival time layer, the distribution of absolute survival differences between contrasting treatment groups resembles that of the pooled population. In this case, using the default $c=20\%$, approximately 20% of non-zero differences between treatment and control pairs will result in ties and move the comparison to the next endpoint (e.g., time to hospitalization). This reduces the influence of random variation on the survival layer. Conversely, when a treatment effect exists on survival time, the distribution of absolute survival differences between treatment groups will shift rightward relative to the pooled distribution. As a result, the same 20% quantile threshold now classifies fewer contrasting pairs as ties in Stage 1, thereby allowing more comparisons to be resolved at the survival layer. Taken together, these properties make the FS-AT test potentially appealing in practice, especially when specifying fixed thresholds a priori is challenging.
@@ 3 Simulation Study
In this section, we examine the performance of the proposed test empirically via simulation. We first compare the proposed test and standard FS test regarding empirical power (after validating their empirical type I error rates), and then investigate the influence that caliper $c$ and weight $w$ have on the empirical power. Although the proposed test can accommodate multiple nonfatal events, we focus on the simpler setting with two endpoints (survival time and time to hospitalization) for a demonstration.
@@ 3.1 Simulation setup
We consider a two-arm clinical trial with a total sample size $n=2000$ and equal allocation, where half of them are assigned to the treatment group $Z=1$ and the other half to the control group $Z=0$. We simulate the two correlated times to represent potentially latent survival time in days and time to hospitalization in days using a copula approach. Following luo2015alternative, we adopt the Gumbel-Hougaard copula with a bivariate distribution that has exponential margins. We let $h_{D}(Z)=\lambda_{D}\exp(-\alpha_{D}Z)$ be the hazard rate for death event, and $h_{T_{1}}(Z)=\lambda_{T_{1}}\exp(-\alpha_{T_{1}}Z)$ be the hazard rate for hospitalization event. Then, the vector of survival time and time to hospitalization (in days) $(D^{*},T_{1}^{*})$ has the joint survival function:
@@
where $\beta\geq 1$ controls the correlation between the two endpoints (i.e., Kendall’s concordance: $1-1/\beta$). We fix parameters $\lambda_{D}=0.0008,\lambda_{T_{1}}=0.0022$ and let $\alpha_{D},\alpha_{T_{1}}\in\{0,0.1,0.2,0.3\}$, in which $\{0,0.1,0.2,0.3\}$ stand for no, very weak, weak, and modest treatment effects. We assume treatment is not worse than control (in terms of extending survival time or time to hospitalization). In our simulation, it is possible to have sample $i$ with $T_{1i}^{*}\geq D^{*}_{i}$, which indicates there is no hospitalization event for $i$. Observed survival time and time to hospitalization are then obtained by performing administrative censoring after the scheduled days of follow-up for all participants, with the scheduled follow-up length denoted as FU. A nominal level of $\alpha=0.05$ for a two-sided test is employed throughout.
We compare the FS-AT test under the default weight and caliper choices (i.e., $\text{FS-AT}(w=1)^{c=20\%}$) with the standard FS test. In addition, we primarily evaluate the performance of the FS-AT test under various choices of weights and calipers, with one exception in assessing the influence of calipers, where a FS-MT test composed of eight stages is included for comparison. Three key factors, follow-up time, correlation between endpoints, and treatment effect size, are varied and detailed in Table 1 for simulation scenarios S1-S8. Additional simulation scenarios are explored in Supporting Information S2. The empirical power is based on 2000 simulation replicates, and the empirical type I error rate calculation is based on 5000 replicates.
All computations are implemented in R, and include function calls from gumbel (gumbel), doParallel (doParallel), dplyr (dplyr), ggplot2 (ggplot2). The reproducible R codes are available in the Supporting Information. An R package “FSMT” that implements the proposed method has been developed and is available at https://github.com/YH-Mou/FSMT.
@@ 3.2 Performance of FS-AT
The empirical power of FS and FS-AT tests under simulation scenarios S1-S8 (i.e., under the non-null scenarios) is summarized in Figure 2. To begin with, in simulation scenarios S1 and S2, the FS-AT test continuously increases the power from the standard FS test by allowing the information on the time-to-hospitalization layer to contribute to the inference about the treatment effect. The difference between FS-AT and FS tests is more apparent with a longer follow-up time. Since longer follow-up time indicates more observed death events on the survival time layer, which brings more random differences, the FS-AT test increases the power by more frequently leveraging the time-to-hospitalization layer and bypassing the random difference on the survival time layer. With a shorter follow-up time, the FS test maintains good performance as lower death rates allow a higher contribution from the time-to-hospitalization layer to pairwise comparison results. In simulation scenarios S3 and S4, the treatment effect is on the survival time layer with random differences on the time-to-hospitalization layer. Although more frequently exploiting the time-to-hospitalization layer is expected to reduce power, the FS-AT test presents comparable power to the FS test in simulation scenario S4, due to employing the proposed adaptive thresholds. In simulation scenario S3, however, it is noticeable that the added correlation between endpoints lowers the power of both methods and the FS-AT test appears to be less powerful than the FS test. This phenomenon will be further investigated in Section 3.3.
@@
Figure 2: Empirical power of FS-AT and FS tests in simulation scenario: (S1) modest treatment effect on the time-to-hospitalization endpoint only, assuming a correlation between survival time and time-to-hospitalization endpoints, (S2) modest treatment effect on the time-to-hospitalization endpoint only, assuming no correlation, (S3) modest treatment effect on the survival time endpoint only, assuming a correlation, (S4) modest treatment effect on the survival time endpoint only, assuming no correlation, (S5) very weak treatment effect on the survival time endpoint and weak treatment effect on the time-to-hospitalization endpoint, assuming a correlation, (S6) very weak treatment effect on the survival time endpoint and weak treatment effect on the time-to-hospitalization endpoint, assuming no correlation, (S7) weak treatment effect on the survival time endpoint and very weak treatment effect on the time-to-hospitalization endpoint, assuming correlation, (S8) weak treatment effect on the survival time endpoint and very weak treatment effect on the time-to-hospitalization endpoint, assuming no correlation.
In simulation scenarios S1-S4, we include only random differences on one of the endpoints to demonstrate the performance of FS-AT test under extreme situations. In practice, however, if prior information indicates no treatment effect on an endpoint, including it in the hierarchical structure should be discouraged when using either FS-AT or standard FS tests. In simulation scenarios S5-S8, we focus on the more common settings where there are treatment effects on both endpoints but with different magnitudes. In simulation scenarios S5 and S6, where the treatment effect on the survival time endpoint is lower than that on the time-to-hospitalization endpoint, the FS-AT test becomes more powerful than the FS test. Similar to simulation scenarios S1 and S2, a longer follow-up time may reduce power as more pairwise comparison results are determined by the survival time endpoint, for which the treatment effect has a weak signal. The FS-AT test appears to be more robust in terms of maintaining power in such a scenario. Different from the trends observed in simulation scenarios S1 and S2, the power of both FS and FS-AT tests here first increases from 250-day follow-up to 500-day follow-up and then starts to decrease, which is likely a result of the difference between having very weak or no treatment effects on the survival time layer. Additionally, in simulation scenarios S7 and S8 where the treatment effect on the time-to-hospitalization endpoint is weak, the FS-AT test achieves comparable power with the FS test (similar to simulation scenario S4). Finally, the empirical type I error in simulation scenario S0 with Kendall’s concordance set to 0 and 0.5 is presented in Table 2. At the nominal level of $\alpha=0.05$, both FS and FS-AT tests adequately control empirical type I error rates, with observed values falling within the expected range of Monte Carlo variation.
@@
Table 2: Empirical type I error rates (%) obtained by FS-AT and FS under the null (simulation scenario S0) with Kendall’s concordance 0 and 0.5. The acceptable range of empirical type I error rate with Monte Carlo variation under 5000 replicates is 4.41% to 5.64% for the 5% nominal level.
Scenario Kendall’s Concordance FU FS FS-AT